# Counting rational points on Hirzebruch–Kleinschmidt varieties over global function fields

### v1

@@ Abstract.
Inspired by Bourqui’s work on anticanonical height zeta functions on Hirzebruch surfaces, we study height zeta functions of split toric varieties with Picard rank 2 over global function fields, with respect to height functions associated with big metrized line bundles. We show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties, where precise analytic properties of the corresponding height zeta functions can be given, allowing for a finer inspection of the asymptotic number of rational points of bounded height on each subvariety.
@@ 1. Introduction
In the papers [FMT89] and [BM90], a program was initiated to understand the asymptotic behavior of the number of rational points of bounded height on Fano algebraic varieties over global fields. In the case of number fields, the current expectation is as follows (see e.g. [Pey03, Formule empirique 5.1] or [ADHL15, Conjecture 6.3.1.5] for precise definitions).
@@
If $\zeta_{U,L}(s)$ converges absolutely in the half-plane $\Re(s)>a>0$, has analytic continuation to $\Re(s)>a-\varepsilon,s\neq a$ for some $\varepsilon>0$ and has a pole of order $b\geq 1$ at $s=a$, then one obtains
@@
(see e.g. [Del54, Théorème III]).
In the case where $K$ is a global function field of positive characteristic, the relevant height functions typically have values contained in $q^{\mathbb{Z}}$, where $q$ the cardinality of the constant subfield $\mathbb{F}_{q}\subset K$. This implies that the associated height zeta functions are invariant under $s\mapsto s+\frac{2\pi i}{\log(q)}$. In particular, having a pole at a point $s=a>0$ implies the existence of infinitely many poles on the line $\Re(s)=a$, which makes it impossible to apply a Tauberian theorem. Moreover, in this setting, it is hopeless to expect an asymptotic formula of the form (3) since $N(U,H_{L},q^{n})=N(U,H_{L},q^{n+\frac{1}{2}})$ would imply $\sqrt{q}=1$ (as explained in [Bou11, Section 1.1]).
For the reasons explained above, for varieties defined over global fields of positive characteristic, the analogue of Conjecture 1.1 is presented in terms of the analytical properties of the respective height zeta function. This is the approach taken by Peyre in [Pey12] in the case of flag varieties. As a variant of Conjecture 1.1, it is then expected that the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$, for appropriate open subsets $U\subseteq X$, converges absolutely for $\Re(s)>1$, it admits meromorphic continuation to $\Re(s)>1-\varepsilon$ for some $\varepsilon>0$, and has a pole at $s=1$ of order $\operatorname{rk}\operatorname{Pic}(X)$ satisfying
@@
and $\tau_{H}(X)$ is the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ defined by Peyre in [Pey12, Section 2] in the case of global fields of positive characteristic. Similarly, as a variant of (2), one expects $\zeta_{U,L}(s)$ to converge absolutely for $\Re(s)>a(L)$ and have meromorphic continuation to $\Re(s)>a(L)-\varepsilon$ for some $\varepsilon>0$ with a pole at $s=a(L)$ of order $b(L)$.
In this paper we focus on smooth projective split toric varieties with Picard rank 2 over a global field $K$ of positive characteristic, and consider height zeta functions associated to big metrized line bundles. Our results show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties for which precise analytic properties of the corresponding height zeta functions can be given. Hence, in this setting, we go beyond the scope of the classical expectations mentioned above. We achieve these results by using concrete defining equations for such varieties and by performing explicit computations on the height zeta functions.
The present work is motivated by the results of Bourqui in [Bou02], [Bou03] and [Bou11], where the above expectation for the anticanonical height zeta function is verified for toric varieties over global fields of positive characteristic, with $U$ the dense torus orbit. Bourqui’s work was inspired by Batyrev and Tschinkel’s proof of Conjecture 1.1 for toric varieties over number fields [BT98].
@@
We refer to these varieties as Hirzebruch–Kleinschmidt varieties. As usual, in order to avoid accumulation of rational points of bounded height, we find it necessary to restrict our attention to rational points in a dense open subset. To this purpose, we define the good open subset $U_{d}(a_{1},\ldots,a_{r})\subset X_{d}(a_{1},\ldots,a_{r})$ as the complement of the closed subvariety define by the equation $x_{tr}=0$.
@@
Let $X=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety of dimension $d=r+t-1$ over a global function field $K=\mathbb{F}_{q}(\mathscr{C})$, let $H=H_{-K_{X}}$ denote the anticanonical height function on $X(K)$, and let $U=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$. Then, the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, and it is a rational function in $q^{-s}$. In particular, it has meromorphic continuation to $\mathbb{C}$. Moreover, it has a pole of order 2 at $s=1$ with
@@
where $|\mathbf{a}|:=\sum_{i=1}^{r}a_{i}$, $h_{K}$ is the class number of $K$, $g$ is the genus of $\mathscr{C}$ and $\zeta_{K}$ is the zeta function of $K$.
@@
Also, since $X$ is a split toric variety, it is well known that $\beta(X)=1$ (see e.g. [BT98, Remark 1.7 and Corollary 1.18]). Hence, Theorem 1.2 and [Bou03, Théorème 1.1] imply that the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ is
@@
As mentioned before, we go beyond the classical expectations and study further the height zeta function of the complement of the good open subset inside $X_{d}(a_{1},\ldots,a_{r})$. This is achieved by a natural decomposition of $X=X_{d}(a_{1},\ldots,a_{r})$ of the form
@@
(see Section 5.4 for details), which allows us to work recursively on the dimension of $X$, and decompose $\zeta_{X,-K_{X}}(s)$ as a finite sum of height zeta functions of projective (and affine) spaces and height zeta functions of good open subsets of Hirzebruch–Kleinschmidt varieties of dimension $d=t+r-1,\ldots,2$. On the one hand, height zeta functions of projective (and affine) spaces are well understood (see Section 4). On the other hand, for each good open subset $U^{\prime}=U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})\subseteq U_{d}(a_{1},\ldots,a_{r})$ with $2\leq t^{\prime}\leq t,1\leq r^{\prime}\leq r$ we get a contribution of a height zeta function of the form $\zeta_{U^{\prime},L}(s)$ with $L\in\operatorname{Pic}(X_{d^{\prime}}(a_{1},\ldots,a_{r^{\prime}}))$ generally distinct from the corresponding anticanonical class. Nevertheless, provided $L$ is big, we can still describe the analytic properties of $\zeta_{U^{\prime},L}(s)$ in detail, see Theorems 5.2 and 5.4 in Section 5.1. Note that the bigness condition on $L$ is necessary, because when $L$ is not big, $\zeta_{U^{\prime},L}(s)$ has no finite abscissa of absolute convergence. The following example illustrates these ideas.
@@
Put $U:=U_{3}(a),U^{\prime}:=U_{2}(a)\subset X^{\prime}:=X_{2}(a)$ and let $\pi^{\prime}:X^{\prime}\to\mathbb{P}^{1}$ denote the corresponding projection map. The anticanonical class on $X$ is $-K_{X}=\mathscr{O}_{X}(2)\otimes\pi^{\ast}(\mathscr{O}_{\mathbb{P}^{2}}(3-a))$. Using Theorem 1.2 we get that $\zeta_{U,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $\mathbb{C}$ with a pole of order 2 at $s=1$ with
$\lim_{s\to 1}(s-1)^{2}\zeta_{U,-K_{X}}(s)=\frac{q^{5(1-g)}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)(a+3)2(q-1)^{2}\log(q)^{2}}.$
By Lemma 5.15 in Section 5.4 the component $U_{2}(a)$ in (6) contributes with the height zeta function $\zeta_{U^{\prime},L}(s)$ with $L:=\mathscr{O}_{X^{\prime}}(2)\otimes(\pi^{\prime})^{\ast}(\mathscr{O}_{\mathbb{P}^{1}}(3-a))\in\operatorname{Pic}(X^{\prime})$, which is big but different from the anticanonical class $-K_{X^{\prime}}$. Nevertheless, by Theorem 5.2 we know that $\zeta_{U^{\prime},L}$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=1$.
Now, by the same Lemma 5.15 the component $\mathbb{P}^{2}$ in (6) contributes with the height zeta function $\zeta_{\mathbb{P}^{2}}((3-a)s)$ where $\zeta_{\mathbb{P}^{2}}(s)$ denotes the “standard height zeta function” of $\mathbb{P}^{2}$ defined in Section 4. We see that $\zeta_{\mathbb{P}^{2}}((3-a)s)$ has no finite abscissa of absolute convergence when $a\geq 3$. Assuming $a<3$, we can apply Theorem 4.1 in Section 4 to deduce that $\zeta_{\mathbb{P}^{2}}((3-a)s)$ converges absolutely for $\Re(s)>\frac{3}{3-a}$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=\frac{3}{3-a}>1$.
Finally, by Lemma 5.15 the component $\mathbb{A}^{1}$ in (6) contributes with the height zeta function $\zeta_{\mathbb{A}^{1}}(2s)=\zeta_{\mathbb{P}^{1}}(2s)-1$, where $\zeta_{\mathbb{P}^{1}}(s)$ denotes the standard height zeta function of $\mathbb{P}^{1}$. In particular, by Theorem 4.1 we know that $\zeta_{\mathbb{A}^{1}}(2s)$ converges absolutely for $\Re(s)>1$ and has meromorphic continuation to $s\in\mathbb{C}$ with a simple pole at $s=1$.
@@
    If $a=1,2$, then the number of rational points of bounded anticanonical height in $X_{3}(a)$ is “dominated” by the number of those points in the component $\mathbb{P}^{2}$ in (6). More precisely, we have that
    $\zeta_{X,-K_{X}}(s)=\zeta_{\mathbb{P}^{2}}((3-a)s)+\zeta_{\mathbb{A}^{1}}(2s)+\zeta_{U,-K_{X}}(s)+\zeta_{U^{\prime},L}(s)$
    converges absolutely for $\Re(s)>\sigma_{a}:=\frac{3}{3-a}>1$ and has meromorphic continuation to $\mathbb{C}$ with a simple pole at $s=\sigma_{a}$, while
    $\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)=\zeta_{\mathbb{A}^{1}}(2s)+\zeta_{U,-K_{X}}(s)+\zeta_{U^{\prime},L}(s)$
    converges absolutely for $\Re(s)>1$. Moreover, by Theorem 4.1 in Section 4, we can compute
    $\lim_{s\to\sigma_{a}}\left(s-\sigma_{a}\right)\zeta_{X,-K_{X}}(s)=\lim_{s\to\sigma_{a}}\left(s-\sigma_{a}\right)\zeta_{\mathbb{P}^{2}}((3-a)s)=\frac{h_{K}q^{3(1-g)}}{(3-a)\zeta_{K}(3)(q-1)\log(q)}.$
@@
    If $a\geq 3$, then the component $\mathbb{P}^{2}$ in (6) has “too many” rational points of bounded height. In fact, $N(\mathbb{P}^{2},-K_{X},B)=\infty$ for all $B\geq 1$. In this case we also have
    $\lim_{s\to 1}\left(s-1\right)^{2}\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)=\lim_{s\to 1}\left(s-1\right)^{2}\zeta_{U,-K_{X}}(s),$
    hence the number of rational points of bounded anticanonical height in $X\setminus\mathbb{P}^{2}$ is “dominated” by the number of those points in the good open subset $U=U_{2}(a)$.
In both cases, we see that only after removing the closed subvariety $\mathbb{P}^{2}\subset X_{3}(a)$ we obtain a height zeta function $\zeta_{X\setminus\mathbb{P}^{2},-K_{X}}(s)$ satisfying the analogue of Conjecture 1.1 over global function fields.
We can also include here the easy case $a=0$ with $X=X_{3}(0)\simeq\mathbb{P}^{1}\times\mathbb{P}^{2}$. Indeed, it follows from Theorem 5.4 in Section 5.1 that $\zeta_{X,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, has meromorphic continuation to $\mathbb{C}$ and has a pole of order 2 at $s=1$ with
$\lim_{s\to 1}(s-1)^{2}\zeta_{X,-K_{X}}(s)=\frac{q^{5(1-g)}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)6(q-1)^{2}\log(q)^{2}}.$
In particular, in this easy case, there is no need to remove a closed subvariety of $X$ in order to verify the analogue of Conjecture 1.1 over global function fields.
@@
The techniques used in this paper are inspired by the work of Bourqui [Bou02] on the anticanonical height zeta function on Hirzebruch surfaces. Combining Bourqui’s ideas with some technical computations, we are able to express the height zeta function $\zeta_{U,L}(s)$, for the good open subset of a Hirzebruch–Kleinschmidt variety and for general big metrized line bundles $L$, essentially as a rational function of degree 2 in the zeta function $\zeta_{K}(s)$ of the base field $K$. In that regard, it would be interesting to investigate if this method can also be applied to other families of algebraic varieties defined over global function fields.
@@
This article is organized as follows. In Section 2 we introduce notation and recall some classical results concerning the geometry of curves defined over finite fields, and analytic properties the zeta functions of global function fields. We also include here a few general properties of toric varieties. In Section 3 we describe the effective cone, anticanonical class, and big divisor classes in Hirzebruch–Kleinschmidt varieties. In Section 4 we recall the analytic properties of the “standard height zeta” function of the projective space. The height zeta functions of Hirzebruch–Kleinschmidt varieties are studied in Section 5. In particular, our main theorems are presented in Section 5.1. Section 5 also includes a precise description of our decomposition of Hirzebruch–Kleinschmidt varieties (see Section 5.4) and the proof of Theorem 1.2 (see Section 5.5). Finally, in Section 6 we apply our general results to Hirzebruch surfaces. In particular, we recover the main result of [Bou02] as a particular case.

### v2

@@ Abstract.
Inspired by Bourqui’s work on anticanonical height zeta functions on Hirzebruch surfaces, we study height zeta functions of split toric varieties with Picard rank 2 over global function fields, with respect to height functions associated with big metrized line bundles. We show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties, where precise analytic properties of the corresponding height zeta functions can be given. As application, we obtain asymptotic formulas for the number of rational points of large height on each subvariety, with explicit leading constants and controlled error terms.
@@ 1. Introduction
In the papers [FMT89] and [BM90], a program was initiated to understand the asymptotic behavior of the number of rational points of bounded height on Fano algebraic varieties over global fields. In the case of number fields, a refined version of the original expectations is as follows (see, e.g., [Pey03, Formule empirique 5.1] or [ADHL15, Conjecture 6.3.1.5] for precise definitions).
@@
Remark 1.2.
In general, the union of all weakly accumulating subvarieties of $X$ is not necessarily contained within a proper closed subset, and more general versions of the Manin–Peyre conjecture suggest that there exists a thin [^2] Following [Ser89], a subset $Z\subset X(K)$ is called thin if it can be expressed as a finite union $Z=\bigcup_{i}\pi_{i}(Y_{i}(K))$ where each $\pi_{i}:Y_{i}\to X$ is a morphism that is generically finite onto its image and has no rational sections. subset $Z\subset X(K)$ such that the asymptotic formula (1) holds for the number of rational points in $X(K)\setminus Z$ of height $\leq B$ (see, e.g., [LT17], [LR19] and references therein). However, for the purpose of this paper, the above formulation is sufficient.
@@
If $\zeta_{U,L}(s)$ converges absolutely in the half-plane $\Re(s)>a>0$, has analytic continuation to $\Re(s)\geq a,s\neq a$ and has a pole of order $b\geq 1$ at $s=a$, then one obtains
@@
(see, e.g., [Del54, Théorème III]).
In the case where $K$ is a global function field of positive characteristic, the relevant height functions typically have values contained in $q^{\mathbb{Z}}$, where $q$ the cardinality of the constant subfield $\mathbb{F}_{q}\subset K$. This implies that the associated height zeta functions are invariant under $s\mapsto s+\frac{2\pi i}{\log(q)}$. In particular, having a pole at a point $s=a>0$ implies the existence of infinitely many poles on the line $\Re(s)=a$, and the strategy described above breaks down. Moreover, in this setting, it is hopeless to expect an asymptotic formula of the form (3) since $N(U,H_{L},q^{n})=N(U,H_{L},q^{n+\frac{1}{2}})$ would imply $\sqrt{q}=1$ (as explained in [Bou11b, Section 1.1]).
For these reasons, for varieties defined over global fields of positive characteristic, the analogue of Conjecture 1.1 is sometimes presented in terms of the analytic properties of the respective height zeta function. This is the approach taken by Peyre in [Pey12] in the case of flag varieties, and by Bourqui in [Bou02], [Bou03] and [Bou11b, Bou16] in the case of Hirzebruch surfaces, split toric varieties, and general toric varieties, respectively. As a variant of Conjecture 1.1, it is then expected that the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$, for appropriate open subsets $U\subseteq X$, converges absolutely for $\Re(s)>1$, it admits meromorphic continuation to $\Re(s)\geq 1$ and has a pole at $s=1$ of order $\operatorname{rk}\operatorname{Pic}(X)$ satisfying
@@
and $\tau_{H}(X)$ is the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ defined by Peyre in [Pey12, Section 2] in the case of global fields of positive characteristic. Similarly, as a variant of (2), one expects $\zeta_{U,L}(s)$ to converge absolutely for $\Re(s)>a(L)$ and have meromorphic continuation to $\Re(s)\geq a(L)$ with a pole at $s=a(L)$ of order $b(L)$.
Remaining in the context of varieties defined over global function fields, there is yet another approach that focuses instead on the asymptotic behaviour of the number
$\tilde{N}(U,H_{L},B):=\#\{P\in U(K):H_{L}(P)=B\}$
of rational points of large height. Asymptotic formulas for these numbers can be deduced, via another Tauberian theorem, from finer analytic properties of the corresponding height zeta functions. This is the approach that we take in this paper.
Remark 1.3.
Given a projective variety $V$ and a smooth projective curve $\mathscr{C}$, both defined over a finite field $\mathbb{F}_{q}$, the study of the number of points in $V(\mathbb{F}_{q}(\mathscr{C}))$ of large height is equivalent to the problem of counting morphisms $\mathscr{C}\to V$ of large degree. This geometric perspective has been pursued by several authors (see, e.g., [Bou11a], [Bou16], [LT19], [Tan21], [BLRT22] and references therein).
In the present paper we focus on smooth projective split toric varieties with Picard rank 2 over a global field $K$ of positive characteristic, and consider height zeta functions associated to big metrized line bundles. Our results show that these varieties can be naturally decomposed into a finite disjoint union of subvarieties for which finer analytic properties of the corresponding height zeta functions can be given. This allows us to obtain asymptotic formulas for the number of rational points of large height on each component of our disjoint union decomposition, going beyond the scope of the classical expectations. Moreover, our asymptotic formulas have controlled error terms and explicit leading coefficients. We achieve these results by using concrete defining equations for our varieties and by performing explicit computations on the height zeta functions.
This work was motivated by the results of Bourqui in [Bou02], [Bou03] and [Bou11b], where the classical expectation for the anticanonical height zeta function is verified for toric varieties over global fields of positive characteristic, with $U$ the dense torus orbit. Bourqui’s work was inspired by Batyrev and Tschinkel’s proof of Conjecture 1.1 for toric varieties over number fields [BT98].
@@
We refer to these varieties as Hirzebruch–Kleinschmidt varieties. As usual, in order to avoid convergence issues, we find it necessary to first restrict our attention to rational points in a dense open subset. To this purpose, we define the good open subset $U_{d}(a_{1},\ldots,a_{r})\subset X_{d}(a_{1},\ldots,a_{r})$ as the complement of the closed subvariety define by the equation $x_{tr}=0$.
@@
Let $X=X_{d}(a_{1},\ldots,a_{r})$ be a Hirzebruch–Kleinschmidt variety of dimension $d=r+t-1$ over the global function field $K=\mathbb{F}_{q}(\mathscr{C})$, let $H=H_{-K_{X}}$ denote the anticanonical height function on $X(K)$, and let $U=U_{d}(a_{1},\ldots,a_{r})$ be the good open subset of $X$. Moreover, put
$\eta_{X}:=\mathrm{g.c.d.}(r+1,t-|\mathbf{a}|).$
Then, the anticanonical height zeta function $\zeta_{U,-K_{X}}(s)$ converges absolutely for $\Re(s)>1$, it is a rational function in $q^{-\eta_{X}s}$ and has a pole of order 2 at $s=1$ with
@@
where $|\mathbf{a}|:=\sum_{i=1}^{r}a_{i}$, $h_{K}$ is the class number of $K$, $g$ is the genus of $\mathscr{C}$ and $\zeta_{K}$ is the zeta function of $K$. Moreover, it is holomorphic for
$\Re(s)>1-\min\left\{\frac{1}{r+1},\frac{1}{(r+1)a_{r}-|\mathbf{a}|+t}\right\},\Re(s)\neq 1,$
and on $\Re(s)=1,s\not\in\left\{1+\frac{2\pi im}{\eta_{X}\log(q)}:m\in\mathbb{Z}\right\}$ it has at most simple poles.
@@
  1. (1)
    The fact that $\zeta_{U,-K_{X}}(s)$ is a function of $q^{-\eta_{X}s}$ follows from the elementary observation that the anticanonical height function $H_{-K_{X}}$ assumes values in $q^{\eta_{X}\mathbb{Z}}$ (see Proposition 3.2(2) and equation (17) in Section 5).
  2. (2)
@@
    Also, since $X$ is a split toric variety, it is well known that $\beta(X)=1$ (see, e.g., [BT98, Remark 1.7 and Corollary 1.18]). Hence, Theorem 1.4 and [Bou03, Théorème 1.1] imply that the Tamagawa number of $X$ with respect to the anticanonical height function $H=H_{-K_{X}}$ is
@@
  3. (3)
    In the case of Hirzebruch surfaces, i.e. $r=1$ and $t=2$, Theorem 1.4 is in accordance with [Bou02, Proposition 3.1] (see also p. 355 in loc. cit.).
As mentioned before, we can deduce from the analytic properties of $\zeta_{U,-K_{X}}(s)$ given in Theorem 1.4, good estimates for the number $\tilde{N}(U,H_{-K_{X}},q^{M})=\#\{P\in K:H_{-K_{X}}(P)=q^{M}\}$ of rational points in $U$ of anticanonical height $q^{M}$ for large $M$. Since this number is trivially zero when $M$ is not divisible by $\eta_{X}$, we can assume $M\in\eta_{X}\mathbb{Z}$.
Corollary 1.6.
Under the hypothesis of Theorem 1.4, for every $\delta>1-\min\left\{\frac{1}{r+1},\frac{1}{(r+1)a_{r}-|\mathbf{a}|+t}\right\}$ we have the estimate
$\tilde{N}(U,H_{-K_{X}},q^{M})=q^{M}(C_{1}M+C_{0}(M))+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\eta_{X}\mathbb{Z},$
where
$C_{1}=\frac{q^{(d+2)(1-g)}\eta_{X}h_{K}^{2}}{\zeta_{K}(t)\zeta_{K}(r+1)((r+1)a_{r}-|\mathbf{a}|+t)(r+1)(q-1)^{2}},$
and $C_{0}(M)\in\mathbb{R}$ depends only on $\frac{M}{\eta_{X}}$ mod $\frac{(r+1)((r+1)a_{r}-|\mathbf{a}|+t)}{\eta_{X}^{2}}$.
We now explain how to study further the number of rational points of large height in the complement of the good open subset of $X_{d}(a_{1},\ldots,a_{r})$. Assume $a_{r}>0$ for simplicity. Then, there is a natural decomposition of $X=X_{d}(a_{1},\ldots,a_{r})$ of the form
@@
(see Section 5.5 for details), which allows us to work recursively on the dimension of $X$, and decompose $\zeta_{X,-K_{X}}(s)$ as a finite sum of height zeta functions of projective (and affine) spaces and height zeta functions of good open subsets of Hirzebruch–Kleinschmidt varieties of dimension $d=t+r-1,\ldots,2$. On the one hand, height zeta functions of projective (and affine) spaces are well understood (see Section 4). On the other hand, for each good open subset $U^{\prime}=U_{t^{\prime}+r^{\prime}-1}(a_{1},\ldots,a_{r^{\prime}})$ with $2\leq t^{\prime}\leq t,1\leq r^{\prime}\leq r$ we get a contribution of a height zeta function of the form $\zeta_{U^{\prime},L}(s)$ with $L\in\operatorname{Pic}(X_{d^{\prime}}(a_{1},\ldots,a_{r^{\prime}}))$ generally distinct from the corresponding anticanonical class. Nevertheless, provided $L$ is big, we can still describe the analytic properties of $\zeta_{U^{\prime},L}(s)$ in detail, see Theorems 5.2 and 5.4 in Section 5.1. Note that the bigness condition on $L$ is necessary, because when $L$ is not big, $\zeta_{U^{\prime},L}(s)$ has no finite abscissa of absolute convergence. The aforementioned results allow us to give precise estimates on the number of rational points of large height on each component in (5); see Theorems 5.16 and 5.19 in Section 5.4. The following example illustrates these results.
@@
Put $U:=U_{3}(a),U^{\prime}:=U_{2}(a)\subset X^{\prime}:=X_{2}(a)$ and let $\pi^{\prime}:X^{\prime}\to\mathbb{P}^{1}$ denote the corresponding projection map. The anticanonical class on $X$ is $-K_{X}=\mathscr{O}_{X}(2)\otimes\pi^{\ast}(\mathscr{O}_{\mathbb{P}^{2}}(3-a))$ and $\eta_{X}=\mathrm{g.c.d.}(2,3-a)$. In particular, $\eta_{X}=1$ if $a$ is even, and $\eta_{X}=2$ if $a$ is odd. Using Corollary 1.6 we get
$\tilde{N}(U,H_{-K_{X}},q^{M})=(C_{1}M+C_{0}(M))q^{M}+O\left(q^{\delta_{1}M}\right)\quad\text{as }M\to\infty,M\in\eta_{X}\mathbb{Z},$
for every $\delta_{1}>\frac{1}{2}$, with $C_{1}>0$ and $C_{0}(M)\in\mathbb{R}$ depending only on the class of $M$ mod $2(a+3)$ if $a$ is even, and only on $\frac{M}{2}$ mod $\frac{a+3}{2}$ if $a$ is odd. By Lemma 5.23 in Section 5.5 the component $U_{2}(a)$ in (6) has height function $H_{L}$ with $L:=\mathscr{O}_{X^{\prime}}(2)\otimes(\pi^{\prime})^{\ast}(\mathscr{O}_{\mathbb{P}^{1}}(3-a))\in\operatorname{Pic}(X^{\prime})$, which is big but different from the anticanonical class $-K_{X^{\prime}}$. Nevertheless, by Theorem 5.16 we get, for every $\delta_{2}>\frac{a+2}{a+3}$, the estimate
$\tilde{N}(U^{\prime},H_{-K_{X}},q^{M})=C_{2}(M)q^{M}+O\left(q^{\delta_{2}M}\right)\quad\text{as }M\to\infty,M\in\eta_{X}\mathbb{Z},$
with $C_{2}(M)>0$ and depending only on the residue class of $M$ mod 2 if $a$ is even, and independent of $M$ if $a$ is odd. Now, by the same Lemma 5.23 the component $\mathbb{P}^{2}$ in (6) has height function $H_{\mathbb{P}^{2}}^{(3-a)}$ where $H_{\mathbb{P}^{2}}$ denotes the “standard height function” on $\mathbb{P}^{2}(K)$ defined in Section 4. We see that $\mathbb{P}^{2}$ has no rational points of height $>1$ when $a\geq 3$. Assuming $a<3$, we can apply the estimate (16) in Section 4 and get
$\tilde{N}(\mathbb{P}^{2},H_{-K_{X}},q^{M})=C_{3}q^{\frac{3M}{3-a}}+O\left(q^{\delta_{3}M}\right)\quad\text{as }M\to\infty,M\in(3-a)\mathbb{Z},$
for every $\delta_{3}>\frac{1}{2(3-a)}$ where $C_{3}>0$, while $\tilde{N}(\mathbb{P}^{2},H_{-K_{X}},q^{M})=0$ if $M\not\in(3-a)\mathbb{Z}$.
Finally, by Lemma 5.23 the component $\mathbb{A}^{1}=\mathbb{P}^{1}\setminus\{[1:0]\}$ in (6) has height function $H_{\mathbb{P}^{1}}^{2}$ where $H_{\mathbb{P}^{1}}$ is the standard height function on $\mathbb{P}^{1}(K)$. In particular, by the estimate (16) we get
$\tilde{N}(\mathbb{A}^{1},H_{-K_{X}},q^{M})=C_{4}q^{M}+O\left(q^{\delta_{4}M}\right)\quad\text{as }M\to\infty,M\in 2\mathbb{Z},$
for every $\delta_{4}>\frac{1}{4}$ where $C_{4}>0$, while $\tilde{N}(\mathbb{A}^{1},H_{-K_{X}},q^{M})=0$ if $M\not\in 2\mathbb{Z}$.
@@
    If $a=1,2$, then $\eta_{X}=3-a=2,1$ respectively, and the number of rational points of large anticanonical height in $X_{3}(a)$ is “dominated” by the number of those points in the component $\mathbb{P}^{2}$ in (6). More precisely, we have that
    $\tilde{N}(X,H_{-K_{X}},q^{M})\sim\tilde{N}(\mathbb{P}^{2},H_{-K_{X}},q^{M})\sim C_{3}q^{\frac{3M}{3-a}}\quad\text{as }M\to\infty,M\in\eta_{X}\mathbb{Z}.$
    Moreover, by (16) we have
    $C_{3}=\frac{h_{K}q^{3(1-g)}}{\zeta_{K}(3)(q-1)}.$
@@
    If $a\geq 3$, then the component $\mathbb{P}^{2}$ in (6) does not contribute to the number of rational points of large height, and $\tilde{N}(X,H_{-K_{X}},q^{M})$ is “dominated” by the number of those points in the good open subset $U=U_{2}(a)$. More precisely, we have
    $\tilde{N}(X,H_{-K_{X}},q^{M})\sim\tilde{N}(U,H_{-K_{X}},q^{M})\sim C_{1}Mq^{M}\quad\text{as }M\to\infty,M\in\eta_{X}\mathbb{Z}.$
    Moreover, we have
    $C_{1}=\frac{q^{5(1-g)}\eta_{X}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)2(a+3)(q-1)^{2}}.$
We can also include here the easier case $a=0$ with $X=X_{3}(0)\simeq\mathbb{P}^{1}\times\mathbb{P}^{2}$. In this case one can directly give a good estimate on the number of rational points of large height in the whole variety $X$. Indeed, a direct application of Theorem 5.19 in Section 5.4 gives, for every $\delta>\frac{1}{4}$, the estimate
$\tilde{N}(X,H_{-K_{X}},q^{M})=(\tilde{C}_{1}M+\tilde{C}_{0}(M))q^{M}+O\left(q^{\delta M}\right)\quad\text{as }M\to\infty,M\in\mathbb{Z},$
with $\tilde{C}_{0}(M)\in\mathbb{R}$ depending only on $M$ mod $6$ and
$\tilde{C}_{1}=\frac{q^{5(1-g)}h_{K}^{2}}{\zeta_{K}(3)\zeta_{K}(2)6(q-1)^{2}}.$
@@
The techniques used in this paper are inspired by the work of Bourqui [Bou02] on the anticanonical height zeta function on Hirzebruch surfaces. Combining Bourqui’s ideas with some technical computations, we are able to express the height zeta function $\zeta_{U,L}(s)$, for $U$ the good open subset of a Hirzebruch–Kleinschmidt variety and for general big metrized line bundles $L$, essentially as a rational function of degree 2 on the zeta function $\zeta_{K}(s)$ of the base field $K$. In that regard, it would be interesting to investigate if this method can also be applied to other families of projective bundles defined over global function fields. We expect to return to this problem in the near future.
@@
This article is organized as follows. In Section 2 we introduce notation and recall some classical results concerning the geometry of curves defined over finite fields, and analytic properties the zeta functions of global function fields. There, we also present the Tauberian theorem that will be applied to our height zeta functions (see Section 2.3), and include a few general properties of toric varieties. In Section 3 we describe the effective cone, anticanonical class, and big line bundle classes in Hirzebruch–Kleinschmidt varieties. In Section 4 we recall an asymptotic formula due to Wan [Wan92] for the number of rational points of large “standard height” on projective spaces over global function fields, and the analytic properties of the associated height zeta functions. The height zeta functions of Hirzebruch–Kleinschmidt varieties are studied in Section 5. In particular, their analytic properties are presented in Section 5.1. These results are used in Section 5.4 to prove our asymptotic formulas for the number of rational points of large height. Section 5 also includes a precise description of our decomposition of Hirzebruch–Kleinschmidt varieties (see Section 5.5) and the proof of Theorem 1.4 (see Section 5.6). Finally, in Section 6 we apply our general results to Hirzebruch surfaces.